Determine the nature of the solutions of the equation.

x^2 - 4x + 6 = 0

The equation has one real solution, two real solutions, or two non real solutions????

Answer by stanbon(75887)

Determine the nature of the solutions of the equation.
x^2 - 4x + 6 = 0
a = 1 ; b = -4 ; c = 6
The equation has one real solution, two real solutions, or two non real solutions????
---
b^2-4ac = 16 - 4*1*6 < 0
---
Ans: 2 non real solutions.
================
Cheers,
Stan H.